Question 9 4 pts Labor demand in a particular market is given by H-800-20W, where H is measured in hours and W is in dollars. In the absence of any government intervention in the labor market, labor supply is H 5W. The equilibrium number of hours worked in this market is H- 160 equilibrium wage in this market is W-$32 round to the nearest integer if necessary) hours. The (enter only numbers in the blanks, and Question 10 4 pts For the labor market described in question 9, calculate the total consumer surplus and the total producer surplus. Remember that in the labor market, firms are the consumers and workers are the producers. The total consumer surplus i10240 The total producer surplus is $ 2560 (enter only numbers in the blanks, and round to the nearest integer if necessary) Hint: It may be helpful to invert the demand and supply functions for the purposes of graphing and calculating consumer and producer surpluses.
Answer #1

9. Equilibrium is where Demand = Supply of labor

800 - 20W = 5W

800 = 25W

W = 800/25

W = 32

H = 5 x 32 = 160 hours

10. H = 800 - 20W

When H = 0 then W = 800/20 = 40

When W = 0 then H = 800

H = 5W

When W = 0 then H = 0

When W = 32 then H = 5 x 32 = 160

Wage Supply 40 C.S 32 P.S Demand 0 160 800 No. of hours worked

Consumer surplus = 1/2 x base x height = 1/2 x 160 x (40 - 32) = 80 x 8 = $ 640

Producer surplus = 1/2 x base x height = 1/2 x 160 x (32 - 0) = 80 x 32 = $ 2560